OpenStudy (anonymous):

simplify this function. sqrt(x^2+1)+sqrt((6-x)^2+4)

OpenStudy (anonymous):

can you do this: y = \[\sqrt(x^2+1)+\sqrt((6-x)^2+4)\] \[y^2 =(x^2+1)+ (6-x)^2+4\] square both sides. y^2 = 2 x^2-12 x+41 y = \[\sqrt{2 x^2-12 x+41}\]

OpenStudy (anonymous):

is this legit?

OpenStudy (anonymous):

Yes, seems to be.

OpenStudy (anonymous):

when i graph it though, via google, they look completely different

hartnn (hartnn):

when you square both sides of \(\large y= \sqrt {a}+\sqrt b\) you DO NOT get y= a+b, you get , y= \(a+b+2\sqrt{ab}\)

Parth (parthkohli):

@dgamma3 Well, to sum up, \((a + b)^2 \ne a^ 2+ b^2\)

Parth (parthkohli):

Yup.\[y^2 = a + b + 2\sqrt{ab}\]

hartnn (hartnn):

because \((m+n)^2=m^2+n^2+\color{red}{2mn}\)

OpenStudy (anonymous):

where am i making this mistake?

Parth (parthkohli):

Going from the first to second step.

OpenStudy (anonymous):

ok, i think i get it

OpenStudy (anonymous):

so, to clarify